John was working at the registration desk of a conference that was scheduled to begin at 8:00 P.M. At 7:30, he texted the conference manager to report that only 20% of the registrants had signed in so far. At 7:45, he sent a second text stating that in the past 15 minutes, 90 additional people had signed in and that now 65% of the participants had arrived. Which of the following is closest to the total number of people who had registered for the conference?
Answer by ikleyn(52781)

.

The problem is to find the unknown x from this equation

    0.65x - 0.2x = 90,


which gives

    0.45x = 90

    x    = 90/0.45 = 200  people.    ANSWER

Solved.